### v1

Abstract
Security conferences are important venues at which academics and practitioners share knowledge about new attacks and state-of-the-art defenses. Despite this, researchers have not studied who shares information and about which security topics. To address this, our study characterizes the speakers, sponsors, and topics presented at the most prestigious academic and industry conferences. We collect a longitudinal data set that contains 9,728 abstracts and 1,686 sponsors across 4 academic and 6 industry conferences. There is limited knowledge sharing between industry and academia. Conferences vary significantly in the equality of how talks/authorship is distributed across individuals. The topics of academic and industry abstracts display consistent coverage of techniques within the MITRE ATT&CK framework. Top tier academic conferences, as well as DEFCON and Black Hat, inconsistently address the governance, response and recovery functions of the NIST Cybersecurity Framework. Commercial InfoSec and insurance conferences (RSA, Gartner, Advisen and NetDillgience) cover the framework more consistently. Prevention and detection remain the most common topic of talks, with no clear temporal trend.
keywords:
cybersecurity, science of security, sociology, information sharing, quantitative social science
1 Introduction
Information security knowledge is produced by both academic and industry researchers. Norms across both communities differ in terms of questions like what “good” research looks like, and the role of peer review However, both communities value gathering in dimly lit conference rooms to listen to researchers explain findings with the support of a slide deck.
To attend conferences, attendees travel to foreign cities, sometimes across continents, and spend thousands of dollars on registration. Although some of these activities are not directly related to formal knowledge exchange (e.g. hallway chats and other extracurricular activities), most participants spend hours listening to talks providing knowledge about novel attacks, mitigation measures, lessons learned, and security concepts. For some attendees in both academia and industry, these conferences are akin to a pilgrimage.
Despite the attendees and their employers investing considerable time and money, security conferences have not been studied as part of the security information sharing ecosystem skopik2016problem ; laube2017strategic ; ainslie2023cyber . We are not aware of research into the speakers or topics at industry conferences. Studies of academic security conferences focus on bibliometrics wendzel2020not ; balzarot2022 and which evaluation method was used in a study carver2016establishing ; burcham2017characterizing .
Further investigation is needed because conferences are not a neutral forum to exchange ideas. While conferences provide a platform for selected researchers to present their work, the low acceptance rates mean other researchers are silenced. In selecting the talks that thousands of attendees watch, program committees control which security topics receive attention. This may influence which solutions are built and implemented. Further, paid speaker slots at industry conferences undermine the so-called market for ideas drezner2017ideas .
These potential problems and the lack of prior research motivate a study that characterizes prestigious security conferences. Our study asks: RQ1: Who speaks at security conferences?; RQ2: About which topics?; and RQ3: How does this vary by conference and over time? We collect a sample of titles, abstracts and speakers/authors across $4$ academic and $6$ industry conferences from 2014 to 2022. To analyze topics, we classified the abstracts into offensive, defensive, or neutral (for theoretical papers). We then mapped offensive talks to the MITRE ATT&CK Framework [^1] https://attack.mitre.org/ (accessed 31 July 2023), and defensive talks to the NIST Cybersecurity Framework [^2] https://www.nist.gov/cyberframework (accessed 31 July 2023).
The rest of this paper is structured as follows. Section 2 discusses related work in the field of information sharing. Section 3 describes our research methodology. Section 4 presents the results of our study. Section 5 discusses the limitations of the study. Section 6 discusses the implications of our results and makes recommendations. Section 7 concludes the paper.
2 Related Work
Defenders use and share many forms of security information to better calibrate their defensive posture to the threat landscape. Most research focuses on sharing real-time technical indicators laube2017strategic , which we review in Section 2.1. Other communication channels exist by which defenders can learn about security, which are covered in Section 2.2.
2.1 Information Sharing
The two most common forms of security information sharing are threat intelligence and vulnerability management. A range of schema and standards have been introduced to facilitate sharing kampanakis2014security , which includes MAEC, OVAL, XCCDF, CPE, CVE, CVSS and so on. Governments have passed regulations to support information sharing yang2020impact , funded Computer Emergency Response Teams to coordinate vulnerability disclosure slayton2020trusting , and directly shared threat intelligence [^3] https://www.cisa.gov/known-exploited-vulnerabilities-catalog. Commercial providers also offer threat intelligence (TI) and surrounding services wagner2019cyber ; sauerwein2017threat ; brown2015cyber ; sauerwein2019analysis . The associated platforms provide information enrichment and analysis functions to obtain targeted and actionable threat intelligence sauerwein2021threat ; dandurand2013towards ; brown2015cyber ; chismon2015threat .
Recent empirical studies have questioned the quality of threat intelligence. Li et al. li2019reading find little overlap between different TI feeds and cannot identify a way to distinguish quality, which confirms findings from other studies thomas2016abuse ; bouwman2022helping . Expert interviews with TI practitioners reveal processes to evaluate quality are informal bouwman2020different , which is to be expected given this remains an open research challenge zibak2022threat . Encouragingly, a study of threat intelligence during COVID-19 found evidence that collaboration, in the form of aggregating commercial TI feeds, improved coverage of threats bouwman2022helping .
Turning to vulnerability management, bug bounties have proliferated, run both independently by vendors zhao2017devising and on platforms like HackerOne sridhar2021hacking . Empirical evaluations support the efficacy of bug bounty programs over internal testing finifter2013empirical ; zhao2015empirical , even though they follow similar processes votipka2018hackers . However, many firms do not patch even when the vulnerability is known and the patch is available frei2006large ; li2017large . This has motivated research into understanding expert users who apply patches li2019keepers ; mathur2016they , and notifications that nudge firms towards patching stock2016hey ; li2016you ; ccetin2019tell .
So much information is shared about security vulnerabilities that prioritizing what to pay attention to has become a research problem jacobs2020improving ; de2023no . Information overload is also a problem for threat intelligence bouwman2020different . This motivates research exploring online advice and knowledge sharing that may be easier for the recipients to digest than technical indicators.
2.2 Advice and Knowledge Sharing
Online forums are often used to solve security and privacy problems li2023itsup ; tahaei2020understanding , including by developers acar2016you . In some cases, this involves copying insecure code snippets directly from the forum post fischer2017stack . This has motivated research into how to nudge developers towards better choices fischer2019stack .
Online security advice targeted at the general public is also widely available redmiles2016think . An analysis of over a thousand sources of advice reveals that the specific recommendations are actionable and comprehensible, albeit with room for improvement redmiles2020comprehensive . However, the sources recommended 374 unique behaviors, which leads to information overload redmiles2020comprehensive . Some users address the quality issue by looking to trusted sources redmiles2016think .
Conferences address this issue by curating the agenda, via peer review in academia. Research into what information is shared at these conferences has been small-scale. For example, Carver et al. carver2016establishing identify which evaluation method (e.g., proof or empirical) was used by each paper in the 2015 IEEE Security and Privacy proceedings. This study was extended to cover the 2016 ACM CCS proceedings burcham2017characterizing . A bibliometric analysis was conducted of around 30,000 academic computer security papers wendzel2020not , focusing on statistics like the number of words, authors, and citations in each paper.
We are not aware of analysis of industry conferences, even though practitioner conferences have tens of thousands of attendees (e.g. 45k at RSA). Across both academia and industry, it remains unclear who speaks/publishes at security conferences (RQ1), about which topics (RQ2), and how this changes over time and conferences (RQ3).
3 Methodology
We focused on the most prestigious conferences because they have the most influence. For each talk, one could analyze article PDFs, video recordings of talks, or meta information like abstracts and titles. Although full academic articles contain the most information, the industry equivalent (white papers) are not available for all talks. Similarly, video recordings of talks are not consistently available, and automated transcription is imperfect. For these reasons, we analyzed abstracts, which are consistently available across industry and academia.
Section 3.1 describes our corpus of conference abstracts. Section 3.2 explains how we used natural language processing to classify what each talk was about, and Section 3.3 describes how we validated this classification. Section 3.4 discusses ethical considerations and open science. Figure 1 provides a high-level overview of our research design.
Figure 1: Our 2-stage classification uses OpenAI’s gpt model in the first step and then both gpt and cosine similarity to map the abstracts to cybersecurity frameworks. We also collect the sponsors of each conference.
3.1 Data Collection
Conferences
In academia, we focused on ‘Top Tier’ computer security conferences soneji2022flawed , namely USENIX Security, IEEE S&P, ACM CCS, and NDSS. We did not consider cryptography venues because they do not map onto either framework (see Section 3.2). Journals were excluded because the lack of in-person presentations undermines the comparison to industry conferences.
For industry, we conducted a web search for articles featuring the ‘Top Security Conferences’. Only Black Hat, DEFCON and RSA were consistently included in all of these articles. For example, the SANS Institute runs influential topic-specific conferences (e.g. Threat Intelligence or Network Security), but no specific event is consistently included as the flagship event. We consider these seven conferences to be the most prestigious computer security conferences.
To diversify our sample, we included Gartner Security because of its influence over procurement decisions and also the two major cyber insurance conferences (Advisen and NetDiligence). Insurance is increasingly used by firms to manage cybersecurity risk dambra2020sok . This makes for an interesting comparison because the insurance community’s roots are in finance rather than IT. Table 1 displays the conferences in our sample, along with some simple statistics.
Conference Sample Window Type Talks per year Characters in title Words per abstract Sponsors Gini Index
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ 1 1/N
Black Hat 2014 - 2022 industry (InfoSec) 252 53.83 54.82 26.43 170.98 99.16 179.0 54.29 0.234 0.381
RSA 2017 - 2022 industry (InfoSec) 338.33 34.98 55.69 17.18 75.75 26.98 75.0a 0.0a 0.288 0.441
DEFCON 2014 - 2022 industry (InfoSec) 92.22 32.19 56.10 25.75 176.72 77.72 0 0 0.179 0.382
Gartner Security 2022 industry (InfoSec) 321.0 0.0 68.07 19.38 66.85 20.38 N/A* N/A* 0.328 0.427
Advisen 2014 - 2022 industry (Insurance) 10.0 4.92 38.81 17.62 58.67 22.25 18.13b 5.95b 0.244 0.383
NetDiligence 2019 - 2022 industry (Insurance) 25.0 6.22 51.50 20.35 36.76 27.36 56.50 15.32 0.159 0.269
USENIX Security 2014 - 2022 academic 132.78 68.33 72.30 22.40 184.12 62.10 15.22 3.33 0.284 0.417
IEEE S&P 2014 - 2022 academic 82.22 34.08 69.77 21.42 222.52 63.53 14.67 6.22 0.280 0.380
ACM CCS 2014 - 2022 academic 166.33 34.47 67.99 21.44 186.21 82.81 13.87c 3.33c 0.325 0.451
NDSS 2014 - 2022 academic 73.0 15.80 72.11 19.79 186.54 82.22 10.14d 3.27d 0.209 0.376
Table 1: Descriptive statistic display the mean value across all years. The Gini index is calculated with the total number of talks per individual across all years. a contains data only from year 2022. b year 2020 is missing. c year 2018 is missing. d only years from 2016 - 2022. N/A* when sponsor data was not available.
Data Extraction
The data was obtained from the conference websites or extracted directly from the proceedings via web content mining. We used a semi-automated process to extract information about speakers, affiliations, talk titles, abstracts, and dates. The academic talks include keynotes and paper presentations for which an abstract was available, but we did not include poster sessions. For industry conferences, we considered all talks with an abstract listed in the conference agenda and ignored sessions such as networking sessions, welcome remarks, and opening speeches.
The data set comprises a total of 9,728 talks (5,639 industry talks and 4,089 academic talks). We also collected 1,686 entries for sponsors of information security conferences, 451 for academic conferences, and 371 for cyber insurance. Some firms sponsored conferences for multiple years.
3.2 Data Analysis
The analysis for RQ1 involved calculating descriptive statistics about the distribution of talks, speakers, and sponsors. Answering RQ2 requires natural language processing (NLP) techniques, as illustrated in Figure 1. We followed a deductive approach by mapping abstracts to established security frameworks, which describe different areas and aspects of defensive and offensive security. The two main decisions were which frameworks and NLP techniques to use.
Choice of Framework The framework would ideally be: (i) comprehensive, (ii) widely used in practice; and (iii) freely available. The first criterion ensures we can meaningfully map more abstracts to the framework, which would not be possible something like a digital forensics standard. The second criterion makes our results easier to understand as some readers will be familiar with the categories from the framework. The final criterion enables others to replicate our work.
For offensive frameworks, MITRE ATT&CK and Cyber Kill Chain framework were considered the most appropriate choices. The ATT&CK framework offers greater depth, satisfying the first criterion. For defensive frameworks, we considered the NIST Cybersecurity Framework, ISO 27001 and CIS Critical Security Controls. We excluded ISO 27001 due to the cost of access. The CIS standard was excluded because NIST CSF is broader, including governance. Although the choice of framework is important, mappings between frameworks are widely available [^4] For example, CIS Controls v8 mapped to NIST CSF is available here: https://www.cisecurity.org/insights/white-papers/cis-controls-v8-mapping-to-nist-csf.
NLP Techniques For the first classification phase, we employed OpenAI’s gpt-3.5-turbo model to classify abstracts into offensive, defensive, or neutral via this prompt:
“is this text cyber security offensive or defensive (just answer one word, namely: offensive, defensive or neutral)? $<$abstract$>$
Neutral talks, which make up a minority (17%) of the corpus, are not further classified.
For the second classification, we referred to the official documentation of each standard to establish themes and sub-themes. For MITRE ATT&CK Framework, we used the 14 tactics as themes and the 196 techniques as sub-themes. For NIST Cybersecurity Framework, we used the 22 categories as themes and the 98 sub-categories as sub-themes. We estimated the similarity between the abstract and the corresponding (sub) theme using two NLP techniques, namely cosine similarity and a Large Language Model.
To calculate cosine similarity, we converted abstracts and each (sub-)theme to a bag of words. We removed stop words using the python nltk stopwords list and then used the nltk wordnet lemmatizer to group the different word forms together. We then compared the talk to each category from the framework using cosine similarity. To compute this, we normalized embeddings with the SentenceTransformer model ‘stsb-mpnet-base-v2’. This process was repeated for the sub-categories of the NIST Cybersecurity Framework and also the techniques from MITRE ATT&CK Framework.
For the Large Language Model, we again used OpenAI’s gpt-3.5-turbo model using the following prompt:
“What are the MITRE ATT&CK tactics that can be identified in the following text. Only return the tactics which are identified: $<$abstract$>$
We used a similar prompt for the categories of NIST CSF. Each theme was individually submitted as a separate query to the model, which produced either a 0 or 1 as output for each talk. To calculate aggregate similarity to each theme, we took the average across all talks in a specific sub-sample.
3.3 Validation
The absence of an existing ground truth mapping for abstract talks to cyber security frameworks made validation difficult. Further, frameworks are high-level with blurry boundaries between categories. To address these issues, we first evaluated the first classification using a hand-labeled data set. Second, we conducted a comparative analysis of the agreement between the two natural language processing techniques. Third, we measured the similarity of offensive, neutral, and defensive talks against both frameworks.
First, we used a human-labeled data set comprising 50 abstracts to evaluate the performance of OpenAI’s gpt-3.5-turbo model at classifying abstracts as offensive, defensive, or neutral. The machine and human agreed on 98% of abstracts.
We then compared the topic classifications across NLP techniques. Since GPT outputs are either 0 or 1, we first apply a threshold to the cosine similarity scores, which are floats between 0 and 1. We set the threshold such that both GPT and cosine similarity would have the same number of 0s and 1s. Otherwise the scores could be perfectly aligned, but appear to disagree purely due to setting different thresholds. For the ATT&CK and NIST Cybersecurity framework categories, the two techniques agreed for 67% and 81% of the abstracts respectively. This ranged across individual categories, from a low of 52% up to 95% (see Table 3). The categories with high accuracy are partly driven by the class imbalance as so few defensive abstracts cover Maintenance (from NIST CSF) or Resource Development (from MITRE ATT&CK).
Third, it might be expected that offensive or defensive talks display the highest similarity to the offensive or defensive frameworks, while neutral talks would exhibit low similarity to both frameworks. Table 2 shows that the cosine similarity scores display the expected ranking. However, this is not true for OpenAI’s gpt-3.5-turbo model, as offensive talks display higher similarity to both frameworks, and defensive talks display the lowest similarity to NIST CSF.
Talk Cosine Similarity gpt-3.5-turbo
Category NIST CSF ATT&CK NIST CSF ATT&CK
Offensive 0.298 0.452 0.160 0.286
Defensive 0.317 0.411 0.135 0.236
Neutral 0.266 0.342 0.149 0.155
Table 2: Average similarity of each category of talks to each framework using both NLP techniques.
Summary Overall, these three validation tests suggest our mapping captures some signal, albeit with considerable noise. Some noise is to be expected given the subjective nature of the task. The GPT method reliably classifies talks as offensive, defensive or neutral. However, the agreement between NLP techniques is disappointing.
Disagreement results from a mixture of errors/hallucinations and also differences in approach. Cosine similarity focuses on keywords and phrases used in the descriptions of each framework. This can lead to false positives when these words are used in a different context, while also leading to false negatives when abstracts use different words with the same meaning. On the other hand, the GPT method captures how words and phrases are used in the data set, which is mainly user-submitted data from the Internet, which is vulnerable to context drift.
Going forward, we will report on aggregate results from OpenAI’s gpt-3.5-turbo because it is a state-of-the-art LLM. Aggregating the classifications helps mitigate random errors as these should cancel out. However, it does not mitigate systemic bias associated with GPT, which we discuss in Section 5. In the Appendix, we also display figures based on an ensemble model, which outputs an equal weighting of the GPT and cosine similarity techniques.
ATT&CK Tactic Agree NIST CSF Function Agree
Credential Access 0.63 Asset Management 0.95
Execution 0.52 Business Environment 0.95
Impact 0.70 Governance 0.90
Persistence 0.56 Risk Assessment 0.84
Privilege Escalation 0.65 Risk Management Strategy 0.93
Lateral Movement 0.79 Supply Chain Risk Mgmt 0.93
Defense Evasion 0.57 Identity Mgmt and Access Control 0.86
Exfiltration 0.71 Awareness and Training 0.90
Discovery 0.52 Data Security 0.90
Collection 0.54 Info. Prot. Processes and Proc. 0.94
Resource Dev. 0.98 Maintenance 0.95
Reconnaissance 0.90 Protective Technology 0.58
Command & Control 0.73 Anomalies and Events 0.65
Initial Access 0.56 Security Continuous Monitoring 0.84
Detection Processes 0.62
Response Planning 0.66
Communications (Respond) 0.78
Analysis 0.67
Mitigation 0.67
Improvements (Respond) 0.65
Recovery Planning 0.79
Improvements (Recover) 0.79
Communications (Recover) 0.81
Mean 0.67 Mean 0.81
Table 3: Fraction of abstracts per category for which the GPT and cosine similarity techniques agree.
3.4 Open Science & Ethics
We will release our corpus of titles, abstracts, speakers, and sponsors in an open source repository, as well as our analysis scripts. We hope this allows other researchers to extend this research project. We discuss these directions in Section 6.
Our main ethical concern is that we process and then share personal data (e.g. author names and affiliations). It would be impracticable to collect affirmative consent from speakers. We believe the benefits in terms of open science outweigh the potential privacy invasion, not least because presenting at conferences is a public act.
4 Results
This section presents the results of our analysis on the composition of talks and sponsors at each conference (Section 4.1), speakers (Section 4.2), and abstract topics (Section 4.3).
4.1 Conference Statistics
RSA and Black Hat had the most talks across the sample period (see Figure 2). Volatility in the size of Black Hat since 2019 is driven by the conference introducing/discontinuing sessions like Day Zero, CISO Summit, and Sponsored Workshops. The academic conferences display the highest rates of growth, as noted by Balzarotti balzarot2022 . USENIX Security doubled in size from 2019 to 2021. The industry conferences do not display the same growth trend. Insurance conferences had the fewest talks, with the absolute number remaining stable over time. This is perhaps surprising given the market grew by over 360% between 2015 and 2021 naic2022report .
Figure 2: Number of talks at each conference over time.
Figure 3: Top sponsors of academic conferences. The number on top of each bar is the mean sponsorship tier of that sponsor.
The academic conferences all had the longest abstracts in terms of word count, which can be considered an imperfect proxy for information content. Table 1 shows that IEEE S&P displays the highest average at 223 words. Abstracts at Black Hat and DEFCON had a similar word count to the other three academic conferences. These five conferences have a mean word count of between 170 and 190, which makes sense given original research is often presented at Black Hat and DEFCON. The conferences with the lowest word counts place less emphasis on research. There was less variance in the length of titles, both within and between conferences.
Black Hat has the most sponsors with an average of 197 per year (see Table 1). It is followed by RSA (75), Netdiligence (57), and Advisen (18). We were unable to find comprehensive sponsorship data for Gartner. The academic conferences all have between 10 and 16 sponsors per year. DEFCON does not accept sponsors. Conferences with fewer sponsors typically had longer abstracts.
Figure 4: Speaker affiliation and sponsor correlation of Black Hat.
The academic conferences are sponsored by a mixture of public and private sources (see Figure 3). Large global technology companies, along with the National Science Foundation, support all four conferences in a similar proportion. InfoSec vendors tend to support specific conferences. Non-profits like the EFF, the FreeBSD Foundation, and DMTF only sponsor USENIX Security.
Although we could not collect information on the price of sponsorship, the numbers on top of the bars in Figure 3 represent the average sponsor tier of that company, where 1 is the highest. This shows that the non-profits like EFF are consistently in lower tiers, while BigTech companies tend to be in the higher tiers. However, the US-based National Science Foundation is most consistently a tier one sponsor. No other country has an equivalent research institution that consistently sponsors these conferences.
The industry conferences are mainly sponsored by InfoSec vendors, with large technology companies sponsoring less frequently. This can be seen in Figure 19 in the Appendix. The analysis of sponsors of industry conferences is less comprehensive because of gaps in the data. The insurance conferences contain a wide range of sponsors (see Figure 19 in the Appendix). These include not just insurance companies, but also law firms, digital forensics providers, communications consultants, and data analytics firms. This is likely because these firms receive work from insurers woods2023lessons .
To probe whether sponsorship influences the number of talks, we analyze the only industry security conference for which we had longitudinal sponsor information, namely Black Hat. Across all years, the conference had $592$ unique sponsors and $1493$ unique affiliation across speakers, which had an intersection of just $219$. This means that 85% of the speakers’ affiliated organizations never sponsored the conference, and 63% of sponsors did not have an affiliated speaker at Black Hat. Figure 4 shows the relationship between number of times an organization sponsored Black Hat and the number of affiliated talks. The two outliers are Microsoft and Google, who respectively had 61 and 55 affiliated talks.
Figure 5: Talks per individual for the top 30 most frequent speakers.
4.2 Speaker Statistics
This subsection examines the distribution of talks among specific individuals. Throughout we assume that each unique name corresponds to a single unique person. Figure 6 shows there is minimal overlap of speakers between industry, academic, and insurance conferences. For this reason, we proceed by analyzing each type of conference separately.
Figure 6: Number of unique speakers at each type of conference.
Figure 5 displays the individuals with the most associations with talks, weighted by $\frac{1}{n}$ where $n$ is the number of authors on a paper or individuals on a panel. Patrick Wardle is a dramatic outlier with 33 talks including more than ten talks at each of Black Hat and DEFCON. The rest of the top 30 industry information security speakers gave between 10 and 15 talks across all four conferences. Some industry speakers presented exclusively at one conference, such as Patrick Hevesi at Gartner. No other Top 30 speaker gave a talk at Gartner. The top insurance speaker had seven talks in total and just 19 speakers gave more than three talks.
Turning to the academic talks, the top authors have over fifty publications, although rarely as a single author. Academics are more likely to publish at multiple conferences than industry speakers. For example, all the top 30 authors have presented at IEEE Security and Privacy, USENIX Security and ACM CCS, although some authors did not present at NDSS. This kind of analysis raises the question of how speaking slots are distributed among participants.
Table 1 displays the Gini index for each conference. A score of 0 means all speakers gave the same number of talks, while a score of 1 means one individual gave all talks. Academic conferences are similarly equal to industry conferences when publications count as $\frac{1}{n}$ where $n$ is the number of co-authors/co-panelists, whereas academic conferences appear considerably less equal when each publications counts as $1$ (see Table 1). This corrects for academics who make small contributions to high $n$ papers—we all have just 24 hours in the day. Both metrics agree that NDSS is the most equal academic conference, while ACM CCS is the least.
Figure 7: Offensive, defensive, and neutral talks at each conference.
4.3 Content of Talks
We now explore the topics of abstracts. Figure 7 shows that defensive talks are more common than offensive and neutral talks at all conferences apart from DEFCON, which is explicitly set up for hackers to present their research. At DEFCON, 64% of the talks were offensive and just 15% (20%) of the talks were defensive (neutral). Across both cyber insurance conferences, just 1% of talks were classified as offensive, whereas 55% of insurance talks were neutral and 44% were defensive. The academic conference all accepted more defensive work than offensive, and some even accepted more neutral than offensive talks.
4.3.1 Defensive Talks
Figure 8 shows that the similarity of talks across areas of NIST CSF has remained relatively stable over time. There is a slight trend of increasing abstracts about risk assessment and governance, as well as decreasing similarity to protective technology. This is generally true when considering trends in topics at individual conferences, and we display some such figures in the Appendix. For this reason, we proceed by analyzing the abstracts aggregated across all years. The NIST Cybersecurity Framework is divided into five core functions: Identify, Protect, Detect, Respond and Recover.
Figure 8: Similarity score of defensive talks mapped to NIST Cybersecurity Framework over years
Before analyzing the aggregate data, there is one exception to the general lack of a temporal trend in topics. Figure 9 shows how DEFCON displayed a considerable temporal shift towards Detection and Response functions of NIST CSF. This could be evidence of the conference’s commercialization, which is an ongoing topic of discussion. The equivalent figures for other conferences can be found in the Appendix.
Figure 9: Similarity score of DEFCON defensive talks mapped to NIST Cybersecurity Framework over years
Identify
The identify function comprises the first six rows in Figure 10. Broadly speaking, Identify covers how organizations make security decisions. Both industry InfoSec and insurance talks address these topics more than the academic talks. The insurance conferences and Gartner display high similarity to Risk Assessment, which captures the likelihood and impact of different security incidents, and Governance, which cover the process of establishing a security policy. It is also notable that the research-oriented industry information security conferences (DEFCON and Black Hat) display less similarity to Identify. Notably, there is low similarity to the final two sections: Risk Management Strategy involves understanding the firm’s meta-reasoning like the assumptions and risk tolerance of the firm, meanwhile Supply Chain Risk Management involves assessing the risk of external firms.
Researchers should not dismiss the Identify function as ‘non-technical’ and outside the scope of computer security. Many of the Identify tasks can be solved using technical data. For example, a recent SoK on asset discovery highlights how Asset Management can be approached using network measurements vermeer2021sok . Similarly, Risk Assessment is analogous to cyber incident prediction, which has been studied as a machine learning problem based on technical indicators and published at top tier venues liu2015cloudy ; bilge2017riskteller .
Figure 10: Similarity score of defensive talks mapped to NIST Cybersecurity Framework.
Protect
The protect function covers the next six rows in Figure 10. Broadly speaking, Protect describes technical controls and organizational processes that directly prevent incidents from occurring. Of these six categories, all conferences had the most similarity to Protective Technology. Figure 11 displays cosine similarity to each sub-category of Protective Technology. The two sub-categories with the highest similarity concern system hardening (PR.PT-3) and network security (PR.PT-4). The next two subcategories concern audits and logging (PR.PT-1), and removable media (PR.PT-2). Finally, resilience mechanisms like fail-safe and load balancing (PR.PT-5) display the lowest similarity.
Academic talks display some similarity to Identity Management and Access Control. This include the evergreen research topic that is passwords morris1979password ; bonneau2012quest . The Awareness and Training category received more coverage in the industry talks relative to the academic talks. Academic research related to this may instead be presented at conferences focusing on Human-Computer Interaction like SOUPS and Chi garfinkel2014usable .
The remaining three categories had low similarity. Maintenance had the lowest similarity across all of the defensive categories. This finding supports the observation that researchers often overlook maintenance even though it is crucial to successful operations collier2020cybercrime . The Information Protection Processes and Procedures category consists of procedures to ensure availability like back-ups, recovery plans and so on. Industry talks more frequently cover these topics. Low similarity to Data Security was perplexing given this is so core to computer security. The first two sub-categories of Data Security, namely data at rest and in transit, are core academic topics. However, the other sub-categories are organizational procedures that tend not to be covered at academic conferences.
Figure 11: Cosine similarity score for the category within Protective Technology.
Detect
The next three rows concern network monitoring and incident detection. Broadly speaking, all of the industry conferences displayed moderate similarity to Security Continuous Monitoring, Anomalies and Events, and Detection Processes. These three categories respectively correspond to: what monitoring technology is in place; how anomalies are identified and responded to; and the surrounding organizational processes like testing, communication, and improvement. NDSS displays higher similarity than the other academic conferences to all three Detect categories, which likely results from the conference’s focus on network security. Across the five NIST Cybersecurity framework functions, the abstracts were most consistently similar to Detect.
Respond
The respond function comprises the next five rows, which are broadly the steps taken to investigate and contain an incident. Abstracts were most similar to Response Planning, which involves establishing incident response procedures before the incident occurs. The talks also displayed reasonable similarity to Mitigation, which involves actions to contain and resolve an incident, and also Improvements, which is about extracting lessons learned from incidents. This is perhaps surprising given the consensus that the security community struggles to extract lessons from incidents knake2021learning .
There was low similarity with Analysis category, which includes forensic investigations and vulnerability disclosure policies, especially at academic conferences. This could be because researchers present at specialist venues like Digital Investigation garfinkel2010digital . There was relatively low similarity to Communications, which involves coordinating with internal and external stakeholders like law enforcement, especially at academic conferences.
Much like with Detect, NDSS displays higher similarity to than the other academic conferences. The insurance conferences, especially NetDilligence, display remarkably high similarity to all five categories. The high level of similarity holds even for Communications despite the other conferences largely ignoring this topic, which supports the theme that insurers have a different perspective on security.
Recover
The final function comprises three categories that all happen post-incident, which all have equivalent categories in the Response function. All conferences show less similarity to the Recover category. This suggests conference talks are more focused on short-term tasks like containing the incident than the long-term goal of restoring the victim to previous function, which is the goal of Recover. The Advisen conference displays the most similarity, which makes sense given recovery is the goal of insurance.
4.3.2 Offensive Talks
We mapped offensive talks to the MITRE ATT&CK framework. This consists of 14 tactics and each tactic has multiple techniques. Gartner, Advisen and NetDilligence are not displayed due to a lack of offensive talks.
Figure 12: Similarity score of offensive talks mapped to MITRE over years
Much like with defensive talks, Figure 12 displays no clear temporal trend in which offensive tactics are covered by conference abstracts. There is, however, a subtle trend towards Impact. We group the 14 tactics into temporal categories: Pre-compromise, First Compromise, Expansion, and Outcome.
Pre-Compromise
The Reconnaissance and Resource Development tactics involve collecting information and other resources (e.g. acquiring exploits) before the attack is launched. Resource development is the tactic that the conferences are least similar to. Reconnaissance, which involves techniques like active scanning and gathering information on the victim’s network, is also very rarely covered by talks. These tactics may be infrequently discussed because they are non-technical.
Figure 13: Similarity score of offensive talks mapped to MITRE ATT&CK
First Compromise
The next four tactics concern how the attacker compromises the first system on a network. Initial Access describes the entry vector into the network, which includes different types of phishing and account compromises. The industry talks conferences display moderate similarity to Initial Access. USENIX Security abstracts are most similar to this among the academic conferences. This breaks from the trend that NDSS talks were most similar to the NIST Cybersecurity Framework out of the academic conferences.
Execution describes tactics that attackers use to run malicious code on the victim’s systems. The industry talks display the most similarity to this category, in particular the Black Hat talks. The academic talks are highly similar to this category, but it is not the highest. Figure 14 displays the similarity of the tactics within Execution.
Persistence describes how attackers maintain access across events like the user rebooting, switching user or any other interruption. It is discussed most often at DEFCON, with Black Hat talks the second most similar. The other conferences display relatively low similarity. This suggests that academic researchers are most concerned by the initial vulnerability, and not how the adversary maintains access to the system.
Privilege Escalation techniques allow an attacker to gain more permissions to access files and other system resources. This technique shows a similar pattern to Persistence, in which academic talks display low similarity relative to industry conferences. USENIX Security talks have the most similarity out of the academic conferences.
Expansion
The next three steps concern how the attacker moves through a victim’s network. Defense Evasion concerns how the adversary avoids detection. The talks display a high similarity to this and also Discovery, which concerns how the adversary collects information about the victim’s network. Unlike the previous tactics, the academic talks display higher similarity to both than the industry talks.
Credential Access involves stealing account names and passwords, such as via keylogging. The conference talks are only moderately similar to this, with academic talks showing more similarity than industry. The conference talks display low similarity to the third tactic in this category, namely Lateral Movement. It covers how the adversary moves through the victim’s network.
Figure 14: Cosine similarity score for the techniques within Execution.
Outcome
These tactics concern what was achieved by the compromise, although these techniques could also be used to expand access. Collection concerns identifying and gathering data on the victim network. It is the tactic in this category with the highest similarity to the conference talks. Meanwhile, Exfiltration concerns how the data is stolen from the network, which has reasonably low similarity to the conference talks.
The Command and Control tactic involves establishing communication channels to maintain control over the infected systems. This has moderate similarities across both industry and academia. The NDSS conference displays the highest similarity, again because this is fundamentally a problem of distributed computing, albeit an illegal one. Finally, the Impact technique concerns an effort to “manipulate, interrupt, or destroy your systems and data”, which would involve ransomware for example. The academic talks display higher similarity to this category, although it is moderate-high for all conferences apart from RSA.
5 Limitations
The four main limitations of our study are: missing data (e.g. sponsorship for some conferences); processing speaker and company names; classifying abstracts against frameworks; and incomplete coverage of security conferences.
Missing Data Our comparative longitudinal study was impacted by changes in reporting and missing data. We could not collect some years of talks due to data unavailability. For example, the RSA conference was created in 1991 but we could only extract the program from 2017 onward. Similarly, there was a lot of missing sponsorship information, which led to a long list of exceptions to the sponsors column in Table 1.
Real Names Counting authors can be complex because two unique names may correspond to the same individual (e.g. Emma Jones and Emma Q. Jones), and one name may correspond to multiple individuals (e.g. Emma Jones from New York and Emma Jones from California). Additionally, some authors may use pseudonyms rather than their real names. For simplicity, we counted unique names as unique people. We considered using affiliations to distinguish between two unique individuals with the same name, but researchers switch institutions too often for this to be reliable.
Similar issues are present for company names. Company names are inconsistently abbreviated (Hewlett Packard vs HP), subsidiaries and even departments are listed (e.g. both Microsoft Corporation and Microsoft Research were listed as sponsors), and some companies change names (e.g. Facebook to Meta). We fixed such issues manually where we could, but we likely missed instances for less well known companies.
Classification Mapping abstracts to categories from security frameworks is challenging because: (i) the task is inherently subjective; (ii) the categories are high-level with fuzzy boundaries; (iii) abstracts typically focus on specific technological systems; and (iv) even state-of-the-art LLMs are known to have errors. These factors help to explain why the agreement between the two NLP techniques (see Section 3.3) was disappointing. Ultimately, we focused on the LLM-based mapping because it is the state-of-the-art technique and it captures how language is used in real discussions in the LLM’s training set. This can be contrasted against the cosine similarity approach, which is trained on a bag-of-words extracted from official documents. However, it is not unreasonable to prefer a different mapping, such as using another NLP technique or another cyber security framework. To enable this, we open-source our data and analysis scripts.
Conference Coverage Our comparative analysis (RQ3) would have been more interesting if we collected data from more conferences and also went further back in time. For example, the FIRST Conference on Computer Security Incident Handling would likely have covered more topics related to Respond, but from a more technical perspective than the insurance conferences. The core challenge to adding more data is that we could only semi-automated our scrapers in order to ensure data quality, and had to create a new parser for each conference. This made it costly to add additional conferences and also to scrape further back in time. Again, we hope that open-sourcing our data and analysis will allow future work to extend our work by adding new conferences.
6 Discussion
We discuss the results and how our framing assumptions shaped the analysis.
Speakers Academia displays more information sharing across conferences than industry, given many industry speakers present at just one conference and speakers rarely cross the insurance–InfoSec line (see Figure 6). It is unclear whether industry conferences would benefit from more cross-pollination. The call for broader information sharing is appealing, however it could be that more efficient communication is enabled by the shared language and assumptions within each community.
A different question is whether sponsorship buys influence drezner2017ideas . Conferences with more sponsors have fewer words per abstract, which could be interpreted as sponsored conferences involving less complex talks. Even though Black Hat had the most sponsors, this appears to have little influence over who speaks. For example, the majority of sponsors (63%) never had an affiliated talk, and the majority (85%) of the speakers’ affiliated organizations never sponsored Black Hat. The main concern for academic conferences is that despite being privacy conferences, two of the five most common sponsors are companies that directly profit from behavioral advertising. The relationship between BigTech companies and researchers has been called into question by other researchers abdalla2021grey .
A similar question is whether it is a problem that a small number of speakers/authors hold the majority of influence. For illustrative purposes, we can compare the Gini index for conferences (see Table 1) to the index of income inequality in specific countries. The distribution of talks across speakers is typically around 0.37–0.44, which is close to the post-tax distribution of income in relatively unequal societies like the US (0.375) and Mexico (0.42). All prestigious industry and academic talks have higher inequality than Germany (0.296). [^5] All figures sourced from the OECD using post-tax income: https://stats.oecd.org Inequality in academic talks is not necessarily a cause for concern. Scientific conferences should provide a platform for the ‘best’ science and that likely means the ‘best’ scientists will present more often. However, an unequal distribution can also result from a narrow perspective of what constitutes ‘good’ research. It is worth noting that ACM CCS was the most unequal academic conference across both approaches to quantifying the Gini Index.
Topics One way to improve inequality metrics is to expand the topics that are presented at top tier academic security conferences, allowing new researchers to present ideas and findings. For example, there is comparatively little research about the Identify function that concerns security management. A first step would be to understand how security is managed at present hielscher2023employees , which can motivate technical solutions to management problems. For example, Asset Management can be inferred from network traffic dekoven2019measuring ; Risk Assessment can be framed as ML prediction liu2015cloudy ; and Governance can involve technical audits rahaman2019security ; mahmud2020cardpliance .
Turning to offensive talks, the academic and industry talks are much more evenly distributed across the techniques in MITRE ATT&CK. The one exception is information gathering that occurs pre-compromise, which is rarely discussed at conferences. Preventing resource development is perhaps better conducted by governments, such as by disrupting criminal markets for exploits allodi2015then . The lack of offensive talks at insurance conferences raises the question of whether there is a sufficient understanding of attacker operations. One potential reason this does not happen is insurers rely on lawyers to manage incidents, and those lawyers do not report findings due to confidentiality concerns woods2023lessons .
Variance The insurance conferences were the biggest outliers because they focus on management and resilience more than technological controls and attacker techniques. This confirms the finding that insurers are more focused on post-incident response than pre-incident mitigation wolff2022cyberinsurance . For the MITRE ATT&CK framework mapping, the only offensive conference (DEFCON) is comparatively focused on how compromise was achieved (execution, persistence and initial access), while the academic conferences focus on activities that take place after compromise.
Turning to changes over time, there is an increasing focus on the Identify function and perhaps a weakening focus on Protective Technology and Detect functions. This contradicts the prediction of Schneier that the 2010s would be the decade of response schneier2014future . Instead, there appears to be a subtle shift to security management. However, we should not overstate these effects as the ‘declining’ categories remain the most frequently discussed in absolute terms.
Assumptions In mapping abstracts to cybersecurity frameworks, we imply that there should be equal coverage across topics of the framework. Our core assumption is that more speakers/authors and different topics is better. We justify this by pointing to the benefits researchers gain from learning about areas outside of their specialism. On the other hand, focusing on ever more specific areas of security facilitates deeper understanding and more progress. This is an example of the explore vs exploit trade-off with an unknown objective function, namely maximizing global security knowledge. We acknowledge that reasonable readers may disagree with our emphasis on explore.
More fundamentally, it is unclear whether encouraging specific lines of work is appropriate, especially given the sources of those frameworks. This has the potential to impinge on academic freedom. Although we explained why ATT&CK and NIST CSF were the best among the alternatives (see Section 3), it is entirely unclear whether they are ‘good enough’ in terms of granularity, completeness and so on. This may be actively contested by some.
Future Work We focused on speakers/authors but did not consider listeners/readers. Doing so is crucial to understanding the impact of conference presentations. Collecting this information risks introducing some form of surveillance that undermines the privacy of attendees. One privacy-preserving approach is to scrape the number of views that uploaded video recordings received and to study the comments. Alternatively, an (imperfect) academic proxy for impact on readers is citation count, but the industry equivalent requires more sophisticated research designs.
To address the limitations described in Section 5, future work should add extra conferences and map abstract to different security frameworks and compare the results. Further, a qualitative study collecting data from conference organizers could explore hidden dynamics, such as negotiations with industry sponsors. This would provide insights into unseen details like what sponsors demand from academic conferences. Gaining access to industry conference organizers is considerably more difficult.
7 Conclusion
Security conferences are an important but understudied channel by which knowledge and information is shared. Our paper built a longitudinal data set of talks at 10 prestigious conferences across industry and academia. We characterized who speaks at these conferences (RQ1), about which topics (RQ2), and how this compares across conferences and over time (RQ3). To answer the latter two research questions, we mapped conference abstracts to cybersecurity frameworks using two NLP techniques, which show moderate agreement.
The results show academics moved freely between the big four venues, but just 5% of academic authors also presented at InfoSec or Insurance conferences. The equivalent figure was 4% for insurance speakers, and 7% for InfoSec speakers, which means there is scope for presenters to move between disciplines. We also showed there is an unequal distribution of talks across speakers, with this inequality varying across conferences. In terms of sponsors, academic conference are dominated by BigTech firms, whereas InfoSec conference sponsors are predominantly cybersecurity vendors.
All conferences contained more offensive than defensive talks, with the exception of DEFCON. Offensive abstracts are distributed evenly across the categories of MITRE ATT&CK, apart from Reconnaissance and Resource Development. Defensive abstracts focused on Protective Technology and categories related to monitoring. Academic conferences tend to overlook the Identify and Recover functions from the NIST Cybersecurity Framework.
These findings support a more nuanced understanding of security information sharing, in which success should be evaluated with regards to specific channels and institutions. Researchers have been disappointed by the quality of information shared via online sources acar2016you ; fischer2017stack and threat intelligence feeds li2019reading ; bouwman2020different . However, there is more success when it comes to vulnerability management. Bug bounties incentivize the discovery of new information finifter2013empirical ; sridhar2021hacking . Further, information about fixes is widely available, the problem lies in acting on the information by installing patches frei2006large ; li2017large . Our results suggest security conferences should be considered a success story, closer to patch management than threat intelligence sharing. This is perhaps not surprising given attendees, sponsors and research founders invest considerable resources in attending and sponsoring these conferences.
Going forward, organizers of prestigious conferences should ensure there is a range of speakers and topics. DEFCON is perhaps a good example in that it refuses sponsorship, evenly covered all but two areas of the MITRE ATT&CK framework, and was one of the most equal conference in terms of the distribution of talks across speakers. An interesting follow-on study would speak to conference organizers to enrich our quantitative analysis with qualitative insights about the unobservable aspects of conference organization.